# A Nonconforming Virtual Element Method for Advection-Diffusion-Reaction Problems with CIP Stabilization

### v1

@@ 3.3 Error Estimates
$u\in H^{2}(\Omega)\cap H^{1}_{g}(\Omega)\cap H^{k+1}(\Omega_{h})\,,$ $f\in H^{k+\frac{1}{2}}(\Omega_{h})\,,$ $\boldsymbol{\beta}\in[W^{k+1}_{\infty}(\Omega_{h})]^{2}\,.$
@@
Some of the other quantities in (97) can be estimated exactly as in [7]. We group them in the following lemma. {comment} QUESTO SOTTO VA RIMOSSO
Lemma 3.8 (Estimates of $\eta_{a}$, $\eta_{c}$, $\eta_{J}$ and $\eta_{\mathcal{F}}$).
Under assumptions (A1) and (A2), for $0<s\leq k$, we have
$\eta_{a}\lesssim\sum_{E\in\Omega_{h}}\varepsilon^{\frac{1}{2}}\,h^{k}\,|u|_{k+1,\mathcal{D}(E)}\,,$ (102)
$\eta_{c}\lesssim\sum_{E\in\Omega_{h}}h^{k+1}\,|u|_{k+1,E}\,,$
$\eta_{J}\lesssim\sum_{E\in\Omega_{h}}h^{s+\frac{1}{2}}\,|u|_{s+1,E}\,,$
$\eta_{\mathcal{F}}\lesssim\sum_{E\in\Omega_{h}}h^{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}k+\frac{1}{2}}}\,|f|_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}k+\frac{1}{2}},E}\,.$
Proof.
VA RIMPIAZZATO CON QUESTI SOTTO
@@
$T^{E}_{b,1}=(\boldsymbol{\beta}\cdot\nabla(u-u_{\mathcal{I}}),v_{h})\lesssim\|\boldsymbol{\beta}\|_{[L^{\infty}(E)]^{2}}|u-u_{\mathcal{I}}|_{1,E}\|v_{h}\|_{0,E}\lesssim h^{k}\|\boldsymbol{\beta}\|_{[L^{\infty}(E)]^{2}}|u|_{k+1,E}\|v_{h}\|_{{\rm cip},E}\,.$
@@
$\lesssim\bigl{(}h^{k}|\boldsymbol{\beta}\cdot\nabla u|_{k,E}+\|\boldsymbol{\beta}\|_{[L^{\infty}(E)]^{2}}|u-u_{\mathcal{I}}|_{1,E}\bigr{)}\|v_{h}\|_{{\rm cip},E}$
$\lesssim h^{k}\|\boldsymbol{\beta}\|_{[W^{k}_{\infty}(E)]^{2}}\|u\|_{k+1,E}\|v_{h}\|_{{\rm cip},E}\,.$
@@
$\lesssim\|\boldsymbol{\beta}\|_{[L^{\infty}(E)]^{2}}|(I-{\Pi^{0,E}_{k}})u_{\mathcal{I}}|_{1,E}\|{\Pi^{0,E}_{k}}v_{h}\|_{0,E}$
$\lesssim\|\boldsymbol{\beta}\|_{[L^{\infty}(E)]^{2}}\bigl{(}|(I-{\Pi^{0,E}_{k}})u|_{1,E}+|(I-{\Pi^{0,E}_{k}})(u-u_{\mathcal{I}})|_{1,E}\bigr{)}\|{\Pi^{0,E}_{k}}v_{h}\|_{0,E}$
$\lesssim h^{k}\|\boldsymbol{\beta}\|_{[L^{\infty}(E)]^{2}}|u|_{k+1,E}\|v_{h}\|_{{\rm cip},E}\,.$
@@
$\lesssim\|\boldsymbol{\beta}\|_{[L^{\infty}(\mathcal{D}(E))]^{2}}\sum_{E^{\prime}\subset\mathcal{D}(E)}\bigl{(}h^{\frac{1}{2}}|u-\Pi^{0}_{k}u|_{1,E^{\prime}}+h^{-\frac{1}{2}}\|u-\Pi^{0}_{k}u\|_{0,E^{\prime}}$
@@
Estimate (105) follows from the bound above, by summing over the elements and taking into account, as already noticed for (82), that each polygon is counted only a uniformly bounded number of times.
@@
$\eta_{\mathcal{N}}=\sup_{v_{h}\in V_{h}^{k,nc}(\Omega_{h})}$ $\Big{[}\frac{1}{\|v_{h}\|_{{\rm cip}}}\sum_{E\in\Omega_{h}}\Big{(}\sum_{e\subset\partial E\cap\Gamma}\varepsilon\langle{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla u_{\mathcal{I}}\cdot\boldsymbol{n}-\nabla u\cdot\boldsymbol{n},v_{h}\rangle_{e}$
$\quad+\sum_{e\subset\partial E\cap\Gamma}\varepsilon\langle u_{\mathcal{I}}-u,{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla v_{h}\cdot\boldsymbol{n}\rangle_{e}$
$\quad+\sum_{e\subset\partial E\cap\Gamma}\dfrac{\varepsilon}{\delta h_{E}}\langle u-u_{\mathcal{I}},\Pi^{0,e}_{k-1}v_{h}\rangle_{e}+\sum_{e\subset\partial E\cap\Gamma_{\text{in}}}\langle|\boldsymbol{\beta}\cdot\boldsymbol{n}|(u-{\Pi^{0,E}_{k}}u_{\mathcal{I}}),{\Pi^{0,E}_{k}}v_{h}\rangle_{e}\Big{)}\Big{]}\,.$
@@
$\sum_{e\subset\partial E\cap\Gamma}\varepsilon\langle$ ${\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla u_{\mathcal{I}}\cdot\boldsymbol{n}-\nabla u\cdot\boldsymbol{n},v_{h}\rangle_{0,e}+\sum_{e\subset\partial E\cap\Gamma}\varepsilon\langle u_{\mathcal{I}}-u,{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla v_{h}\cdot\boldsymbol{n}\rangle_{0,e}$ (107)
$\quad+\dfrac{\varepsilon}{\delta h_{E}}\sum_{e\subset\partial E\cap\Gamma}\langle u-u_{\mathcal{I}},\Pi^{0,e}_{k-1}v_{h}\rangle_{e}+\sum_{e\subset\partial E\cap\Gamma_{\text{in}}}\langle|\boldsymbol{\beta}\cdot\boldsymbol{n}|(u-{\Pi^{0,E}_{k}}u_{\mathcal{I}}),{\Pi^{0,E}_{k}}v_{h}\rangle_{0,e}$
@@
For the first term we have
$T_{N,1}^{E}$ $=\sum_{e\subset\partial E\cap\Gamma}\varepsilon\langle\nabla u\cdot\boldsymbol{n}^{E}-\nabla{\Pi^{\nabla,E}_{k}}u_{\mathcal{I}}\cdot\boldsymbol{n}^{E},v_{h}\rangle_{e}$ (108)
$\lesssim\sum_{e\subset\partial E\cap\Gamma}\varepsilon\|\nabla u-\nabla{\Pi^{\nabla,E}_{k}}u_{\mathcal{I}}\|_{0,e}\|v_{h}\|_{0,e}$
$\lesssim\varepsilon h^{-\frac{1}{2}}\left(\|\nabla u-\nabla{\Pi^{\nabla,E}_{k}}u_{\mathcal{I}}\|_{0,E}+h\,|\nabla u-\nabla{\Pi^{\nabla,E}_{k}}u_{\mathcal{I}}|_{1,E}\right)\sum_{e\subset\partial E\cap\Gamma}\|v_{h}\|_{0,e}$
@@
$\lesssim\varepsilon^{\frac{1}{2}}h^{k}|u|_{k+1,E}\|v_{h}\|_{{\rm cip},E}\,.$
For the second one we get
$T_{N,2}^{E}$ $=\sum_{e\subset\partial E\cap\Gamma}\varepsilon\langle u_{\mathcal{I}}-u,{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla v_{h}\cdot\boldsymbol{n}\rangle_{0,e}$ (109)
$\lesssim\sum_{e\subset\partial E\cap\Gamma}\varepsilon\|u-u_{\mathcal{I}}\|_{0,e}\|{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla v_{h}\|_{0,e}$
@@
$\lesssim\left(\dfrac{\varepsilon}{\delta h_{E}}\sum_{e\subset\partial E\cap\Gamma}\|u-u_{\mathcal{I}}\|_{0,e}^{2}\right)^{\frac{1}{2}}\left(\dfrac{\varepsilon}{\delta h_{E}}\sum_{e\subset\partial E\cap\Gamma}\|\Pi^{0,e}_{k-1}v_{h}\|_{0,e}^{2}\right)^{\frac{1}{2}}$
@@
$\lesssim\varepsilon^{\frac{1}{2}}h^{k}|u|_{k+1,E}\|v_{h}\|_{{\rm cip},E}$
For the last term we get
$T_{N,4}^{E}$ $=\sum_{e\subset\partial E\cap\Gamma_{\text{in}}}\langle|\boldsymbol{\beta}\cdot\boldsymbol{n}|(u-{\Pi^{0,E}_{k}}u_{\mathcal{I}}),{\Pi^{0,E}_{k}}v_{h}\rangle_{e}$ (111)
$\lesssim\sum_{e\subset\partial E\cap\Gamma_{\text{in}}}\|u-{\Pi^{0,E}_{k}}u_{\mathcal{I}}\|_{0,e}\,\||\boldsymbol{\beta}\cdot\boldsymbol{n}|^{\frac{1}{2}}{\Pi^{0,E}_{k}}v_{h}\|_{0,e}$
@@
{comment}
Lemma 3.13 (Estimate of $\eta_{J}$).
Under the assumptions (A1) and (A2), the term $\eta_{J}$ can be bounded as follows (for $0<s\leq k$)
$\eta_{J}\lesssim h^{s+\frac{1}{2}}\,|u|_{s+1,E}\,.$
Combining Lemmas 3.6-3.12 with Proposition 3.2, we obtain the following result.
@@
with constants $\Theta_{E}$ depending on $|u|_{{k+1},E}$, $|f|_{{k+\frac{1}{2}},E}$, and $\|\beta\|_{[W^{k+1}_{\infty}(E)]^{2}}$, but independent of $h$ and $\varepsilon$.
@@ 4 Numerical results
In this section we numerically test our method, by considering two model problems in the domain $\Omega=[0,1]\times[0,1]$. We use two different families of meshes:
@@
$+\sum_{e\in\mathcal{E}_{h}^{\partial}{,\ e\subset\Gamma_{\text{in}}}}\left\||\boldsymbol{\beta}\cdot\boldsymbol{n}|^{{\frac{1}{2}}}\Pi^{0,E}_{k}(u-u_{h})\right\|^{2}_{0,e}+J(u-u_{h},u-u_{h}).$
@@
In Figure 3, we observe the results for $u_{1}$ and $u_{2}$, when the method orders are $k=1,2,3$. Optimal rates of convergence in the $L^{2}-$norm and $H^{1}-$seminorm can be appreciated. For the CIP-norm and $k=1$, we observe a super-linear convergence rate. This behavior can be explained by considering the error estimate of Theorem 3.3: most likely, for such errors and small $\varepsilon$, the dominating part is the second one in the right-hand side of (112), which correspond to a convergence rate of order $\frac{3}{2}$, since $s=k=1$.
@@
As expected, in Figure 4, we observe that the CIP norm of the error is almost constant. Left column is related to $u_{1}$, right column is related to $u_{2}$. Similar results not reported here have been obtained using other meshes.

### v2

@@ 3.3 Error Estimates
$u\in H^{2}(\Omega)\cap H^{1}_{g}(\Omega)\cap H^{k+1}(\Omega_{h})\,,$ $f\in H^{k+\frac{1}{2}}(\Omega_{h})\,,$ $\boldsymbol{\beta}\in[W^{k+1}_{\infty}(\Omega_{h})]^{d}\,.$
@@
Some of the other quantities in (86) can be estimated exactly as in [7]. We group them in the following lemma.
@@
$T^{E}_{b,1}=(\boldsymbol{\beta}\cdot\nabla(u-u_{\mathcal{I}}),v_{h})\lesssim\|\boldsymbol{\beta}\|_{[L^{\infty}(E)]^{d}}|u-u_{\mathcal{I}}|_{1,E}\|v_{h}\|_{0,E}\lesssim h^{k}\|\boldsymbol{\beta}\|_{[L^{\infty}(E)]^{d}}|u|_{k+1,E}\|v_{h}\|_{{\rm cip},E}\,.$
@@
$\lesssim\bigl{(}h^{k}|\boldsymbol{\beta}\cdot\nabla u|_{k,E}+\|\boldsymbol{\beta}\|_{[L^{\infty}(E)]^{d}}|u-u_{\mathcal{I}}|_{1,E}\bigr{)}\|v_{h}\|_{{\rm cip},E}$
$\lesssim h^{k}\|\boldsymbol{\beta}\|_{[W^{k,\infty(E)}]^{d}}\|u\|_{k+1,E}\|v_{h}\|_{{\rm cip},E}\,.$
@@
$\lesssim\|\boldsymbol{\beta}\|_{[L^{\infty}(E)]^{d}}|(I-{\Pi^{0,E}_{k}})u_{\mathcal{I}}|_{1,E}\|{\Pi^{0,E}_{k}}v_{h}\|_{0,E}$
$\lesssim\|\boldsymbol{\beta}\|_{[L^{\infty}(E)]^{d}}\bigl{(}|(I-{\Pi^{0,E}_{k}})u|_{1,E}+|(I-{\Pi^{0,E}_{k}})(u-u_{\mathcal{I}})|_{1,E}\bigr{)}\|{\Pi^{0,E}_{k}}v_{h}\|_{0,E}$
$\lesssim h^{k}\|\boldsymbol{\beta}\|_{[L^{\infty}(E)]^{d}}|u|_{k+1,E}\|v_{h}\|_{{\rm cip},E}\,.$
@@
$\lesssim\|\boldsymbol{\beta}\|_{[L^{\infty}(\mathcal{D}(E))]^{d}}\sum_{E^{\prime}\subset\mathcal{D}(E)}\bigl{(}h^{\frac{1}{2}}|u-\Pi^{0}_{k}u|_{1,E^{\prime}}+h^{-\frac{1}{2}}\|u-\Pi^{0}_{k}u\|_{0,E^{\prime}}$
@@
Estimate (93) follows from the bound above, by summing over the elements and taking into account, as already noticed for (77), that each element is counted only a uniformly bounded number of times. ∎
@@
$\eta_{\mathcal{N}}=$ $\sup_{v_{h}\in V_{h}^{k,nc}(\Omega_{h})}\Big{[}\frac{1}{\|v_{h}\|_{{\rm cip}}}\sum_{E\in\Omega_{h}}\Big{(}\sum_{e\subset\partial E\cap\Gamma}\varepsilon\langle{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla u_{\mathcal{I}}\cdot\boldsymbol{n}^{E}-\nabla u\cdot\boldsymbol{n}^{E},v_{h}\rangle_{e}$
$\quad+\sum_{e\subset\partial E\cap\Gamma}\varepsilon\langle u_{\mathcal{I}}-u,{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla v_{h}\cdot\boldsymbol{n}^{E}\rangle_{e}$
$\quad+\dfrac{\varepsilon}{\delta h_{E}}\sum_{e\subset\partial E\cap\Gamma}\langle u-u_{\mathcal{I}},\Pi^{0,e}_{k-1}v_{h}\rangle_{e}+\sum_{e\subset\partial E\cap\Gamma_{\text{in}}}\langle|\boldsymbol{\beta}\cdot\boldsymbol{n}^{E}|(u-{\Pi^{0,E}_{k}}u_{\mathcal{I}}),{\Pi^{0,E}_{k}}v_{h}\rangle_{e}\Big{)}\Big{]}\,.$
@@
$\sum_{e\subset\partial E\cap\Gamma}\varepsilon\langle$ ${\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla u_{\mathcal{I}}\cdot\boldsymbol{n}^{E}-\nabla u\cdot\boldsymbol{n}^{E},v_{h}\rangle_{e}+\sum_{e\subset\partial E\cap\Gamma}\varepsilon\langle u_{\mathcal{I}}-u,{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla v_{h}\cdot\boldsymbol{n}^{E}\rangle_{e}$ (95)
$\quad+\dfrac{\varepsilon}{\delta h_{E}}\sum_{e\subset\partial E\cap\Gamma}\langle u-u_{\mathcal{I}},\Pi^{0,e}_{k-1}v_{h}\rangle_{e}+\sum_{e\subset\partial E\cap\Gamma_{\text{in}}}\langle|\boldsymbol{\beta}\cdot\boldsymbol{n}^{E}|(u-{\Pi^{0,E}_{k}}u_{\mathcal{I}}),{\Pi^{0,E}_{k}}v_{h}\rangle_{e}$
@@
For $T_{N,1}^{E}$, we have
$T_{N,1}^{E}$ $=\sum_{e\subset\partial E\cap\Gamma}\varepsilon\langle{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla u_{\mathcal{I}}\cdot\boldsymbol{n}^{E}-\nabla u\cdot\boldsymbol{n}^{E},v_{h}\rangle_{e}$ (96)
$\leq\sum_{e\subset\partial E\cap\Gamma}\varepsilon\|\nabla u-{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla u_{\mathcal{I}}\|_{0,e}\|v_{h}\|_{0,e}$
$\lesssim\varepsilon h^{-\frac{1}{2}}\left(\|\nabla u-{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla u_{\mathcal{I}}\|_{0,E}+h\,|\nabla u-{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla u_{\mathcal{I}}|_{1,E}\right)\sum_{e\subset\partial E\cap\Gamma}\|v_{h}\|_{0,e}\,.$
Since it holds (cf. Lemma 2.1 and Lemma 2.3):
$\|\nabla u-{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla u_{\mathcal{I}}\|_{0,E}$ $\leq\|\nabla u-{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla u\|_{0,E}+\|{\boldsymbol{\Pi}^{0,E}_{k-1}}(\nabla u-\nabla u_{\mathcal{I}})\|_{0,E}$
$\lesssim h^{k}|u|_{k+1,E}\,,$
and
$|\nabla u-{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla u_{\mathcal{I}}|_{1,E}$ $\leq|\nabla u-{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla u|_{1,E}+|{\boldsymbol{\Pi}^{0,E}_{k-1}}(\nabla u-\nabla u_{\mathcal{I}})|_{1,E}$
$\lesssim|\nabla u-{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla u|_{1,E}+h^{-1}\|{\boldsymbol{\Pi}^{0,E}_{k-1}}(\nabla u-\nabla u_{\mathcal{I}})\|_{0,E}$
$\lesssim h^{k-1}|u|_{k+1,E}\,,$
we get
$T_{N,1}^{E}$ $\lesssim\varepsilon h^{k}|u|_{k+1,E}\,h^{-\frac{1}{2}}\sum_{e\subset\partial E\cap\Gamma}\|v_{h}\|_{0,e}$ (97)
@@
$\lesssim\varepsilon^{\frac{1}{2}}h^{k}|u|_{k+1,E}\|v_{h}\|_{{\rm cip},E}\,,$
where, in the penultimate step, we have used the inequality established in the proof of Lemma 2.2.
For $T_{N,2}^{E}$, we get
$T_{N,2}^{E}$ $=\sum_{e\subset\partial E\cap\Gamma}\varepsilon\langle u_{\mathcal{I}}-u,{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla v_{h}\cdot\boldsymbol{n}^{E}\rangle_{e}$ (98)
$\leq\sum_{e\subset\partial E\cap\Gamma}\varepsilon\|u-u_{\mathcal{I}}\|_{0,e}\|{\boldsymbol{\Pi}^{0,E}_{k-1}}\nabla v_{h}\|_{0,e}$
@@
$\leq\left(\dfrac{\varepsilon}{\delta h_{E}}\sum_{e\subset\partial E\cap\Gamma}\|u-u_{\mathcal{I}}\|_{0,e}^{2}\right)^{\frac{1}{2}}\left(\dfrac{\varepsilon}{\delta h_{E}}\sum_{e\subset\partial E\cap\Gamma}\|\Pi^{0,e}_{k-1}v_{h}\|_{0,e}^{2}\right)^{\frac{1}{2}}$
@@
$\lesssim\varepsilon^{\frac{1}{2}}h^{k}|u|_{k+1,E}\|v_{h}\|_{{\rm cip},E}.$
Finally, for $T_{N,4}^{E}$, we get
$T_{N,4}^{E}$ $=\sum_{e\subset\partial E\cap\Gamma_{\text{in}}}\langle|\boldsymbol{\beta}\cdot\boldsymbol{n}^{E}|(u-{\Pi^{0,E}_{k}}u_{\mathcal{I}}),{\Pi^{0,E}_{k}}v_{h}\rangle_{e}$ (100)
$\lesssim\sum_{e\subset\partial E\cap\Gamma_{\text{in}}}\|u-{\Pi^{0,E}_{k}}u_{\mathcal{I}}\|_{0,e}\,\||\boldsymbol{\beta}\cdot\boldsymbol{n}^{E}|^{\frac{1}{2}}{\Pi^{0,E}_{k}}v_{h}\|_{0,e}$
@@
Combining Lemmas 3.63.11 with Proposition 3.2, we obtain the following result.
@@
with constants $\Theta_{E}$ depending on $|u|_{k+1,E}$, $|f|_{k+\frac{1}{2},E}$, and $\|\beta\|_{[W^{k+1,\infty}(E)]^{2}}$, but independent of $h$ and $\varepsilon$.
@@ 4 Numerical results
In this section we numerically test our method in two space dimensions, by considering two model problems in the domain $\Omega=[0,1]\times[0,1]$. We use two different families of meshes:
@@
$+\sum_{e\in\mathcal{E}_{h}^{\partial},\ e\subset\Gamma_{\text{in}}}\left\||\boldsymbol{\beta}\cdot\boldsymbol{n}^{E}|^{\frac{1}{2}}\Pi^{0,E}_{k}(u-u_{h})\right\|^{2}_{0,e}+J(u-u_{h},u-u_{h}).$
@@
In Figure 3, we observe the results for $u_{1}$ and $u_{2}$, when the method orders are $k=1,2,3$. Optimal rates of convergence in the $L^{2}-$norm and $H^{1}-$seminorm can be appreciated. For the CIP-norm and $k=1$, we observe a super-linear convergence rate. This behavior can be explained by considering the error estimate of Theorem 3.3: most likely, for such errors and small $\varepsilon$, the dominating part is the second one in the right-hand side of (101), which correspond to a convergence rate of order $\frac{3}{2}$.
@@
As expected, in Figure 4, we observe that the CIP norm of the error is almost constant. Left column is related to $u_{1}$, right column is related to $u_{2}$. Similar result not reported here have been obtained using other meshes.